OpenStudy (anonymous):

Express the following equation in slope-intercept form: y + 7 = -3/4 (x - 20)

OpenStudy (amistre64):

- out the 7 and multiply your slope thru the (...) and simply as wanted

OpenStudy (amistre64):

y + 7 = -3/4 (x - 20) - 7 -7 --------------------- y = -(3/4)x -(-3(-20)/4) - 7

OpenStudy (amistre64):

that -20 shoulda been just a "20"

OpenStudy (amistre64):

y = -(3/4)x -(-3(20)/4) - 7 y = -(3/4)x -(-3(5)) - 7 y = -(3/4)x -(-15) - 7 y = -(3/4)x +15 - 7 y = -(3/4)x +8 if i kept it straight